You are considering purchasing one of two assets. • Asset 1 has payments of 5,000 at the end of year 1, 10,000 at the end ofA 2.26% B 2.79% 3.71% D 5.45% E 4.73%
Answer #1

OPTION D

Asset 1:
26000=5000/1.03+10000/(1+s3)^3+15000/1.0425^5
=>(1+s3)^3=10000/(26000-5000/1.03-15000/1.0425^5)
=>(1+s3)^3=1.11559259892

Asset 2:
29500=12000/(1+1 year rate deferred 3 years)+20000/(1.0425^5/(1+s3)^3)
=>1 year forward rate deferred 3 years=12000/(29500-20000/(1.0425^5)*1.11559259892)-1=5.4470%